OpenStudy (anonymous):

express the following as a single logarithm ln(8+x^6)+(1/2)lnx-ln(cosx)

OpenStudy (anonymous):

\[1/2\ln x=\ln x^{1/2}\] \[\ln(a)+\ln(b)-\ln(c)=\ln \left(\frac{ ab }{ c }\right)\]

OpenStudy (anonymous):

ln[(x^(1/2)*(8+x^6))/(cosx)] would this be the correct answer?

OpenStudy (anonymous):

yes